A male patient has the following respiratory values. What is the patient's calculated total lung capacity (TLC)?

Tidal volume (TV) = 512 ml

Expiratory reserve volume (ERV) = 1245 ml

Vital capacity (VC) = 4732 ml
Answer #1

Tidal volume (TV) ,= 512 ml  

Expiratory reserve volume (ERV) = 1245 ml

Vital capacity (VC) = 4732 ml

total volume of lungs is sum of vital capacity & residual volume: TLC = VC + RV+TV

=4372 + 1245 + 512 =6129 ml

the patient's calculated total lung capacity=6129 ml